# Liu, Rahimian, and Yu

### v5

@@ 2.2 Influence Sample Differential Privacy
For any $\epsilon$-node DP seeding algorithm $M^{k}(G)$ with $n>1$ and $\epsilon>0$, there exist graphs $G$ on $n$ nodes for which the following holds. When the maximum degree of $G$ is n1n-1, $\max_{S:|S|=1}I_{G}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\frac{n}{4}-(e^{\epsilon}+1)$. When the maximum degree satisfies $2\leq d<n-1$, $(1-1/e)\max_{S:|S|=1}I_{G}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\frac{1}{10e^{\epsilon}}\Bigl(1-\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)\Bigr)n-1+\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)$.
When the maximum degree of $G$ is n1n-1, we use the star graph, since it differs from a totally isolated graph by the connection of a single node; a node-DP algorithm must output similar seed sets on these two graphs and thus yields low-quality seeds. When the maximum degree satisfies $2\leq d<n-1$, we construct $G^{\prime}=(V,p^{\prime})$ on $n$ nodes consisting of $l$ disjoint d1d-1-regular trees $S_{1},\dots,S_{l}$, each with roughly $n/l$ nodes and edges $p^{\prime}_{e}=1$ within trees and $p^{\prime}_{e}=0$ otherwise, and obtain $G=(V,p)$ by adding one node $v$ that connects the root nodes of two trees $S_{i}$ and $S_{j}$. Under node DP, the algorithm must produce outputs with similar probabilities on $G$ and $G^{\prime}$, so it cannot assign substantially higher probability to the high-quality seeds in $G$ than in $G^{\prime}$, where such seeds are ineffective, leading to low expected influence. Based on these constructions, the following proposition (proved in Section 6.2) further shows that no edge-DP seeding algorithm can output high-quality seeds either, and both hardness results extend to the approximate $(\epsilon,\delta)$-DP setting where Equation 1 allows an additive $\delta$; see Section 6.2.
@@
For any $\epsilon$-edge DP seeding algorithm $M^{k}(G)$ with $n>1$ and $\epsilon>0$, there exists a graph $G$ on $n$ nodes whose maximum degree satisfies $d\geq 2$ such that $(1-1/e)\max_{S:|S|=1}I_{G}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\frac{1}{10e^{\epsilon}}\Bigl(\frac{4}{5}-\frac{2}{e}\Bigr)n-\frac{4}{5}+\frac{2}{e}.$.
@@ 6 Proofs and Supplementary Results
For graphs with maximum degree $2\leq d<n-1$, we construct $G^{\prime}=(V,p^{\prime})$ consisting of $l$ disjoint (d1)(d\!-\!1)-regular trees $S_{1},\ldots,S_{l}$, each of size about $n/l$ (a complete d1d-1-regular tree truncated to the largest depth not exceeding $n/l$, with the last level partially filled if necessary). For each edge $e$, set $p_{e}^{\prime}=1$ if $e$ belongs to some $S_{i}$ and $p_{e}^{\prime}=0$ otherwise. By the pigeonhole principle, there exist $a$ trees $S_{i_{1}},\ldots,S_{i_{a}}$ such that $\Pr[M^{k=1}(G^{\prime})\subset S_{i_{1}}\cup\cdots\cup S_{i_{a}}]\leq a/l$.
We then obtain $G=(V,p)$ by adding a new node $v$ and connecting it to the roots of $S_{i_{1}},\ldots,S_{i_{a}}$. Any $u\in S_{i_{1}}\cup\cdots\cup S_{i_{a}}$ now reaches all nodes in these $a$ trees, giving $I_{G}(u)=an/l+1$, while any $u\notin S_{i_{1}}\cup\cdots\cup S_{i_{a}}$ reaches only its own tree, so $I_{G}(u)=n/l$.
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\leq\frac{ae^{\epsilon}}{l}\Bigl(\frac{an}{l}+1\Bigr)+\Bigl(1-\frac{ae^{\epsilon}}{l}\Bigr)\frac{n}{l}=\Bigl(\frac{1}{a}+\frac{(a-1)e^{\epsilon}}{l}\Bigr)\frac{an}{l}+\frac{ae^{\epsilon}}{l}.$
Comparing with $\mbox{\sf OPT}=an/l$ gives
$(1-1/e)\mbox{\sf OPT}-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\Bigl(\frac{a-1}{a}-\frac{1}{e}-\frac{(a-1)e^{\epsilon}}{l}\Bigr)\frac{an}{l}-\frac{ae^{\epsilon}}{l}.$
@@
$(1-1/e)\mbox{\sf OPT}-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M(G)}[I_{G}({\mathbf{y}})]\geq\frac{1}{10e^{\epsilon}}\Bigl(1-\frac{d}{d-1}\Bigr)\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)n-1+\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr).$

### v6

@@ 2.2 Influence Sample Differential Privacy
For any $\epsilon>0,n>1$, and $\epsilon$-node DP seeding algorithm $M^{k}$, there exists a graph $G$ of $n$ nodes so that $\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\leq\mbox{\sf OPT}-\left(\frac{n}{4}-(e^{\epsilon}+1)\right)$ where $\mbox{\sf OPT}=\max_{S:|S|=1}I_{G}(S)$.
Moreover, for any $2\leq d$, there exists a graph $G$ of $n$ nodes and maximum degree upper bounded by $d$ so that $\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\leq(1-1/e)\mbox{\sf OPT}-\frac{1}{10e^{\epsilon}}\Bigl(1-\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)\Bigr)n+\Bigl(1-\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)\Bigr)\Bigl(1+\frac{1}{10e^{\epsilon}}\Bigr)$.
When the maximum degree of $G$ is n1n-1, we use the star graph, since it differs from a totally isolated graph by the connection of a single node; a node-DP algorithm must output similar seed sets on these two graphs and thus yields low-quality seeds. When the maximum degree satisfies $2\leq d<n-1$, we construct $G^{\prime}=(V,p^{\prime})$ on $n$ nodes consisting of $l$ disjoint d1d-1-regular trees $S_{1},\dots,S_{l}$, each with roughly (n1)/l(n-1)/l nodes and edges $p^{\prime}_{e}=1$ within trees and $p^{\prime}_{e}=0$ otherwise, and obtain $G=(V,p)$ by adding one node $v$ that connects the root nodes of two trees $S_{i}$ and $S_{j}$. Under node DP, the algorithm must produce outputs with similar probabilities on $G$ and $G^{\prime}$, so it cannot assign substantially higher probability to the high-quality seeds in $G$ than in $G^{\prime}$, where such seeds are ineffective, leading to low expected influence. Based on these constructions, the following proposition (proved in Section 6.2) further shows that no edge-DP seeding algorithm can output high-quality seeds either, and both hardness results extend to the approximate $(\epsilon,\delta)$-DP setting where Equation 1 allows an additive $\delta$; see Section 6.2.
@@
For any $n>1$, $\epsilon>0$, and $\epsilon$-edge DP seeding algorithm $M^{k}$, there exists a graph $G$ on $n$ nodes whose maximum degree satisfies $d\geq 2$ such that $\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\leq(1-1/e)\mbox{\sf OPT}-\frac{1}{10e^{\epsilon}}\Bigl(\frac{4}{5}-\frac{2}{e}\Bigr)n+\frac{4}{5}-\frac{2}{e}$ where $\mbox{\sf OPT}=\max_{S:|S|=1}I_{G}(S)$.
@@ 6 Proofs and Supplementary Results
For graphs with maximum degree $2\leq d<n-1$, we construct $G^{\prime}=(V,p^{\prime})$ consisting of $l$ disjoint (d1)(d\!-\!1)-regular trees $S_{1},\ldots,S_{l}$, each of size about (n1)/l(n-1)/l (a complete d1d-1-regular tree truncated to the largest depth not exceeding (n1)/l(n-1)/l, with the last level partially filled if necessary). For each edge $e$, set $p_{e}^{\prime}=1$ if $e$ belongs to some $S_{i}$ and $p_{e}^{\prime}=0$ otherwise. By the pigeonhole principle, there exist $a$ trees $S_{i_{1}},\ldots,S_{i_{a}}$ such that $\Pr[M^{k=1}(G^{\prime})\subset S_{i_{1}}\cup\cdots\cup S_{i_{a}}]\leq a/l$.
We then obtain $G=(V,p)$ by adding a new node $v$ and connecting it to the roots of $S_{i_{1}},\ldots,S_{i_{a}}$. Any $u\in S_{i_{1}}\cup\cdots\cup S_{i_{a}}$ now reaches all nodes in these $a$ trees, giving $I_{G}(u)=a(n-1)/l+1$, while any $u\notin S_{i_{1}}\cup\cdots\cup S_{i_{a}}$ reaches only its own tree, so $I_{G}(u)=(n-1)/l$.
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\leq\frac{ae^{\epsilon}}{l}\Bigl(\frac{a(n-1)}{l}+1\Bigr)+\Bigl(1-\frac{ae^{\epsilon}}{l}\Bigr)\frac{n-1}{l}=\Bigl(\frac{1}{a}+\frac{(a-1)e^{\epsilon}}{l}\Bigr)\frac{a(n-1)}{l}+\frac{ae^{\epsilon}}{l}.$
Comparing with $\mbox{\sf OPT}=a(n-1)/l$ gives
$(1-1/e)\mbox{\sf OPT}-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\Bigl(\frac{a-1}{a}-\frac{1}{e}-\frac{(a-1)e^{\epsilon}}{l}\Bigr)\frac{a(n-1)}{l}-\frac{ae^{\epsilon}}{l}.$
@@
$(1-1/e)\mbox{\sf OPT}-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M(G)}[I_{G}({\mathbf{y}})]\geq\frac{1}{10e^{\epsilon}}\Bigl(1-\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)\Bigr)n-\Bigl(1-\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)\Bigr)\Bigl(1+\frac{1}{10e^{\epsilon}}\Bigr).$